If n is an integer, which of the following CANNOT be a factor of 3n + 4 ?
(A) 4 (B) 5 (C) 6 (D) 7 (E) 8

Answer by MRperkins(300)

Set 3n+4 = to each possible solution and then solve for n.
3n+4=4
3n+4=5
3n+4=6
3n+4=7
3n+4=8
solve each one and if it is an integer(...,-2,-1,0,1,2,3,...) then it will work.
3n+4=4
3n=0
n=0 (0 is an integer and therefore (A)4 is a solution
.
3n+4=5
3n=1
n=1/3 (1/3 is not an integer and therefore (B)5 is NOT a solution.
.
continue with the others. There are 2 solutions: A and ?